Suppose a ball is drawn at random from a box containing three white and three black balls. After a ball is drawn, it is then replaced and another drawn. What is the probability that of the first four balls drawn, exactly two are white?
1
Expert's answer
2020-06-02T20:18:21-0400

According to the binomial distribution formula the probability of drawn exactly 2 white balls is:

"P(2)=C_4^2\cdot(\frac{1}{2})^2\cdot(\frac{1}{2})^2=0.375"